A fair coin is tossed until a head appears.The probability that 3 throws are necessary.
I considered 1/2*1/2*1/2=1/8
But the answer is 1/2*1/2 DISABLED_event_only= 1/4.
Please suggest.

Answer by ewatrrr(24785)

 
Hi
consider the sample space:
P(3 throws necessary) = 1-6/8 = 2/8 = 1/4
THT
THH
HHH
HHT
HTH
HTT
TTT
TTH